Flying Cloud Co. has the following operating data for its manufacturing operations:

Unit selling price $236
Unit variable cost $105
Total fixed costs $799,000

The company has decided to increase the wages of hourly workers which will increase the unit variable cost by 10%. Increases in the salaries of factory supervisors and property taxes for the factory will increase fixed costs by 4%. If sales prices are held constant, the next break-even point for Flying Cloud Co. will be

a.increased by 637 units

b.increased by 797 units

c.decreased by 797 units

d.increased by 956 units
Answer #1

Current Breakeven point = Fixed cost/Contribution margin per unit

= 799,000/(236-105) = 6099 units

Proposed breakeven point

= (799,000+4%) /[(236 - (105*110%)]

= 830,960/120.50

= 6896 units.

Break-even point increased by 797 units (6896 - 6099)

Option B is the answer